OpenStudy (anonymous):

The ratio of the length (L) to width (w) of a certain building is equal to the ratio of the height (h) to girth (g) of a man. Which equation represents the man’s height?

OpenStudy (anonymous):

I believe it is h= g(l/w) because: l/w = h/g solve for h (l/w)*g = h